OpenStudy (anonymous):

At what points does the helix r(t) = sin t, cos t, t intersect the sphere x^2 + y^2 + z^2 = 82? (Round your answers to three decimal places. If an answer does not exist, enter DNE.) (x, y, z) = __________________(smaller t-value) (x, y, z) = ____________________larger t-value

OpenStudy (anonymous):

Hmmm, maybe solve for \(t\) when \[ (\sin(t))^2 + (\cos(t))^2+(t)^2 = 82 \]

ganeshie8 (ganeshie8):

where were u stuck at ?

ganeshie8 (ganeshie8):

\((\sin(t))^2 + (\cos(t))^2+(t)^2 = 82\) \(\sin^2(t) + \cos^2(t)+t^2 = 82\)

ganeshie8 (ganeshie8):

u need to use a trig identity : \(\sin^2 \theta + \cos^2\theta = 1\)

ganeshie8 (ganeshie8):

\(\sin^2(t) + \cos^2(t)+t^2 = 82\) \(1+t^2 = 82 \) solve \(t \)

OpenStudy (anonymous):

for smaller t-value ( -0.909,-0.416,-2)???

ganeshie8 (ganeshie8):

how did u get t = -2 ?

OpenStudy (anonymous):

-9:S

ganeshie8 (ganeshie8):

then ?

ganeshie8 (ganeshie8):

r(t) = <sin t, cos t, t > r(-9) = ?